Ten (10) identical cars are going to compete in a race. Reliability of each car for the duration of one race is 0.95. What is the probability that at least one of these race cars fails during the race?
Answer #1

Ans:

Reliability or probability of not failing a car during the race=0.95

There are 10 identical cars.

Probability that at least one of these cars fails during the race=1-P(none of the cars fails)

=1-0.95^10

=1-0.5987

=0.4013